Prove by mathematical induction that n^2+n < 2^n whenever n is an integer greater

than 4.


1
Expert's answer
2021-11-22T06:19:53-0500

"\text{\underline{Inducive base:} We check for $n=5$, then we have the following: } \\
5^2 + 5 = 30 < 32 = 2^5 \\
\text{which shows that the statement is true for } n=5.\\
\text{\underline{Inducive hypothesis:} we now assume that the statement is true for every $k > 4$ i.e} \\
k^2 + k < 2^k , k<4\\
\text{\underline{Inductive step:}}\\
(k+1)^2+k+1\\
=k^2+k+2k+2 < 2^k+2^k + 2\quad -(*)\\
\text{Note that $2^k= \displaystyle{\sum^k_{r=0}\begin{pmatrix}k\\r \end{pmatrix}}$} \\
\implies 2^{k+1}= 2\cdot\displaystyle{\sum^k_{r=0}\begin{pmatrix}k\\r \end{pmatrix}}\\
\implies \text{for k > 4, }2^k>2n+2\\
\text{From * we have that}\\
(k+1)^2+k+1<2^k+2^k\\
=2\cdot 2^k = 2^{k+1}\\
\text{Hence we have our result}"